# High-speed ditching of double curvature specimens with cavitation and ventilation

### v1

@@ 2.5 Test Conditions
The test conditions that are discussed in the present paper are listed in Table 2. All tests presented here are performed guide inclination angle $\beta=\arctan(V/U)$. Being $V/U=1.5/40=0.0375$, then $\beta=2.1476^{\circ}$. It is worth noticing that, due to the use of bungee ropes to accelerate the trolley, a very precise control of the speed cannot be achieved, which is the reason why the actual values of the horizontal speed are not exactly the nominal expected values. It should be noted that in the discussion from now on, however, reference will be made to the nominal velocity values and not to the actual ones reported in the table.
@@ 3 Results and discussion
3.1 Definition of the impact phase and of $\tau$
For the purpose of the present study, the data analysis is limited to the impact phase, which lasts from the time of first contact of the specimen with the water surface, $t_{0}$, up to the time $t_{E}$ at which the root of the forward propagating spray reaches the leading edge of the specimen. As done in Iafrati & Grizzi (2019), $t_{0}$ can be approximated as the time at which the first pressure probe exhibits a pressure growth, whereas $t_{E}$ is estimated by linear extrapolation starting from the times at which the pressure peak passes over the location of the probes P28 and P30 (Iafrati, 2016). As shown in Figure 4, for the shapes S2 and S3 at a pitch angle of 6 the estimation of $t_{0}$ based on the time history of the pressure measured by the probe P17 is expected to be quite accurate, being the probe P17 located very close to the point of first contact H. Instead, for the shape S1B the first contact point is located between P21 and P24, therefore the estimate of $t_{0}$ is more uncertain. The estimate based on pressure measurements is corrected by computing the time duration of the impact phase from the underwater videos and subtracting it from $t_{E}$.
@@ 3.2 Hydrodynamics phenomena during the water entry
Qualitative considerations on the hydrodynamics of the water entry of the double curvature specimens can be made through a comparative analysis of the pressure time histories and the underwater high-speed videos. Such an analysis is conducted on the test of the shape S2, advancing with a horizontal velocity $U$=40 m/s, velocity ratio $V/U$=0.0375 and pitch angle $\alpha$= 6. Video frames taken at different $\tau$, displaying the flow evolution on the bottom of the specimen, are shown in Figure 7.
@@
Figure 7: Frames of the underwater video at different $\tau$ during the water entry of the shape S2 at a nominal horizontal speed of $U$=40 m/s and at a pitch angle of $\alpha$=6. The specimen moves from left to right.
@@
(a) Probe positions: the point C separates the part of the specimen with zero curvature on the right, from that with finite curvature on the left; H is the point of first contact at $\alpha$=6.
@@
Figure 8: Pressure probe time histories row by row for the shape S2 moving at $U$=40 m/s, at a pitch angle of $\alpha$=6.
From the underwater frames shown in Figure 7 it is possible to recognize the wetted area as the darker region. During the early stage of impact, i.e. in Figure 7$(b,c,d)$, the wetted area grows in all directions. The brighter zone around the wetted area represents the spray region.
@@
By focusing on the longitudinal plane (midplane) in the trolley-fixed reference frame, the results of the two-dimensional case of an asymmetrical wedge entering water with a horizontal velocity (Judge et al., 2004; Krastev et al., 2018) can provide some elements for the interpretation of the flow phenomena in the present case. In fact, as a consequence of the kinematic asymmetry of the flow due to the horizontal speed, the pressure on the front part of the specimen is higher than that at the rear. From a theoretical standpoint, it is expected that a velocity singularity will occur at the apex (Judge et al., 2004). In real flows, a strong local velocity acceleration is observed there, accompanied by a pressure drop, which is similar to what occurs on the suction side of hydrofoils (Faltinsen, 2005; Newman, 1971). The flow field presents also similarities with that of the vertical water entry of inclined asymmetric wedges (Semenov & Iafrati, 2006; Riccardi & Iafrati, 2004). In Xu et al. (2008), in which a wedge entering water with both a vertical and a horizontal velocity is studied using a self-similar approach, even if the velocity ratio $U/V$ is much lower than the present case, a local pressure drop to negative values is observed at the wedge apex.
An alternative interpretation of the pressure behaviour can be obtained by focusing on the flow in the transverse plane, in an earth-fixed reference frame, by exploiting the so called 2D+$t$ approach (Fontaine & Tulin, 1998; Iafrati & Broglia, 2010). In the front part the plate always penetrates into the water, whereas in the rear part, depending on the local inclination of the tangent to the body surface in the longitudinal plane and on the $U/V$ velocity ratio, the fluid flow problem changes from a water entry to a water exit one (Tassin et al., 2013) and when this happens, negative pressures develop. The development of low and even negative pressures occurs very shortly after the first contact with water and, based on the above considerations, the pressure drop is expected to be higher if the gradient in the slope of the body surface is more significant, which means in presence of larger longitudinal curvature.
@@
As observed in Iafrati & Grizzi (2019), the pressure drop can be so high that the vapour pressure value may be reached. Given the water temperature of the tests between 15 and 18  C, the vapour pressure relative to the atmospheric value, denoted as $p_{\textrm{vap}}$, is in the range -99.6 and-99.2 kPa. This explain the formation of a cavitation region (Acosta, 1973; Brennen, 2014), as shown in the underwater images of Figure 7 at the non-dimensional time $\tau\geq 0.2$. In the trolley-fixed reference frame, the cavitation region has a sort of reverse ”D” shape, which follows the pressure distribution on the bottom of the specimen. As the specimen enters in water, the forward front of the cavitation region maintains rather straight and located about the point of curvature change C, whereas the rear edge of the cavitation region spreads backwards towards the trailing edge of the specimen. Once the cavitation region reaches the trailing edge of the specimen, the water vapour inside it, at vapour pressure, comes into contact with air at atmospheric pressure. As a consequence, air is suddenly entrained into the water vapour cavity, forming a ventilation zone that expands in the opposite direction, i.e towards the leading edge. This behaviour is typical of the so called cavitation-ventilation regime, described in more detail in Section 3.3 based on the analysis of the water entry of the specimen S3 (Iafrati & Grizzi, 2019; Iafrati et al., 2020) and also occurs for S2 at the same horizontal speed.
@@
The expansion of the cavitation region can be also recognized by observing the time histories of the pressure probes, see Figure 8. The passage of the rear edge of the cavitation region corresponds to the abrupt drops to the vapour pressure value. Moving backwards, i.e. from row $r4$ to row $r1$, the instant at which the drops occur for the midline probes (P17, P13, P9 and P4) increases. This time delay is related to the propagation velocity of the rear edge of the cavitation region. Comparing the pressure time histories across the same rows, it can be noted that pressure drops of the midline probes are delayed with respect to those of the probes at the sides, due to the combined effect of of the body shape and of the curvature of the rear edge of the cavitation region. When ventilation starts and air is entrained within the cavity, pressures start increasing progressively, reaching the zero-pressure value. It is worth noting that, in contrast to the pressure drop associated with the propagation of the cavitation region, the rise in pressure values during this phase is gradual and occurs at extremely close time instants for all pressure probes positioned on the same row. The increase in pressure values occurs later in time, when progressing from row $r1$ to $r4$. This clearly indicates that the ventilation front is advancing at an extremely high speed, as previously observed for the shape S3 in Iafrati & Grizzi (2019).
@@
A comprehensive global overview of the shape and evolution of the cavitation region can be derived by constructing pressure iso-lines using the available pressure data from row $r1$ to $r5$ at different time instants. In Figure 9 the iso-lines overlapped to the corresponding video frames at $\tau$=0.3, 0.6 and 0.8 are shown.
(a) $\tau$= 0.3
(b) $\tau$= 0.6
(c) $\tau$= 0.8
@@
Figure 9: Pressure iso-lines in the rear part of the specimen overlapped to the underwater video frames for the test of shape S2 at U=40 m/s, at $\alpha=6^{\circ}$ and at $\tau$=0.3, 0.6 and 0.8. The symmetry about the mid-line is exploited.
@@
Observing the pressure time histories, it can be observed that ventilation occurs within the time interval between $\tau\approx 0.7$ and 0.8. By focusing the pressure iso-lines over this time interval, as shown in Figure 10, it is possible to gain more insight on the ventilation dynamics.
(a) Shape S2 - $\tau$=0.70
(b) Shape S2 - $\tau$=0.72
(c) Shape S2 - $\tau$=0.74
(d) Shape S2 - $\tau$=0.76
(e) Shape S2 - $\tau$=0.78
(f) Shape S2 - $\tau$=0.8
[…]
In fact, it is observed that the pressure begins to rise back to the atmospheric value first at the centre, since the cavitation region reaches the trailing edge first there. However, the ventilation front develops rapidly and at about $\tau$=0.78. becomes roughly parallel to the transverse direction, as also observed for the shape S3 in Iafrati & Grizzi (2019).
Given the considerations made in this section, based on the observation of the underwater video-frames and on the pressure time histories, a schematic showing the evolution of the cavitation region and the formation of the ventilation zone in the mid-plane is proposed in Figure 11.
Figure 11: Qualitative schematic describing the evolution of the cavitation region and of the ventilation zone in the mid-plane.
@@
Figure 12: Time histories of the forces recorded by the front (red) and rear (blue) load cells and of the total normal load (black), for shape S2 at 6 pitch angle and $U$=40 m/s. The vertical red and blue lines indicate the start of cavitation and of ventilation respectively. The solid and dashed lines refer to the first and second test repeat respectively (denoted as R1 and R2).
@@
3.3 Effect of the horizontal speed
@@
As previously noted in Iafrati & Grizzi (2019); Iafrati et al. (2020), significant alterations in the hydrodynamics of the water entry are observed when the horizontal speed is varied for the shape S3. In particular, at least four regimes were identified: a no-cavitation regime (for $U$ lower than about 27 m/s), in which no cavitation is observed; an incipient cavitation regime (between $U\approx$ 27 m/s and 35 m/s) in which a cavitation region expanding in time only up to a certain $\tau$ can be seen; an intermediate regime, (between $U\approx$ 35 m/s and 37 m/s) (denoted in Iafrati & Grizzi (2019) as Full Cavitation Bubble and Cavitation Alternate with Ventilation), in which the cavitation region reaches the trailing edge, followed by the formation of a local ventilation region, which does not expand in relevant way; finally a cavitation-ventilation regime ($U$ higher than 37 m/s), in which the ventilation region exhibits a successive fast expansion towards the leading edge of the specimen.
@@
In the present work, the effect of the horizontal speed is analysed for the shape S2, entering water with a pitch angle of 6 and at velocity ratio of $V/U=0.0375$. The analysis is performed in terms of pressures and loads and through the observation of the underwater images taken at two non-dimensional times, i.e. $\tau=0.2$ and $\tau=0.6$. A comparison is presented between the time histories of pressures recorded by the probes located on the midline and on line $l3$. (Figure 6). In order to make the comparison easier, the probes lying along rows $r5-r8$ and $r1-r4$ are plotted in separate graphs. The results for $U$=30 m/s and 35 m/s are shown in Figure 13, whereas those for $U$=40 m/s and 45 m/s in Figure 14.
(a) $U$=30 m/s - Frame $\tau$ =0.2
(b) $U$=35 m/s - Frame $\tau$ =0.2
(c) $U$=30 m/s - Frame $\tau$ =0.6
(d) $U$=35 m/s - Frame $\tau$ =0.6
(e) $p$ front probes - midline - $U$= 30 m/s
(f) $p$ front - probes midline - $U$= 35 m/s
(g) $p$ front probes - line l3 $U$= 30 m/s
(h) $p$ front probes - line l3 $U$= 35 m/s
(i) $p$ midline - rear probes - $U$= 30 m/s
(j) $p$ midline - rear probes - $U$= 35 m/s
(k) $p$ line l3 $U$= 30 m/s
(l) $p$ line l3 $U$= 35 m/s
@@
Figure 13: Front and rear pressure probe time histories and under-water video frames comparison for the tests of shape S2, at a pitch angle 6 and $U$=30 m/s and 35 m/s.
(a) $U$=40 m/s - Frame $\tau$ =0.2
(b) $U$=45 m/s - Frame $\tau$ =0.2
(c) $U$=40 m/s - Frame $\tau$ =0.6
(d) $U$=45 m/s - Frame $\tau$ =0.6
(e) $p$ front probes - midline - $U$= 40 m/s
(f) $p$ front - probes midline - $U$= 45 m/s
(g) $p$ front probes - line l3 $U$=40 m/s
(h) $p$ front probes - line l3 $U$=45 m/s
(i) $p$ midline - rear probes - $U$=40 m/s
(j) $p$ midline - rear probes - $U$=45 m/s
(k) $p$ line l3 $U$=40 m/s
(l) $p$ line l3 $U$= 45 m/s
@@
Figure 14: Front and rear pressure probe time histories and under-water video frames comparison for the tests of shape S2, at a pitch angle 6 and $U$=40 m/s and 45 m/s.
@@
As shown in Figures 13 and 14 the probes located in the front, i.e. ahead of the point H, are characterized by the typical sharp rise, followed by a gentle decay. It is worth noting that, for all probes, the non-dimensional time $\tau$ at which the pressure peak is attained is almost independent of the horizontal speed, whereas the peak value increases with the speed. As expected, the peaks recorded by the probes along the line $l3$ occur at a later time with respect to the peaks recorded by the probes at the midline in the same row. This is a consequence of the combined effect of the transverse curvature of the specimen and of the curvature of the spray root, which is associated with the possibility for the water to escape from the sides (Iafrati, 2016).
@@
The analysis of the pressure measurements of the probes located behind the point H is less straightforward and the support of the underwater images is needed for a correct interpretation. In fact, looking at the underwater frames in Figures 13 and 14 it can be seen that at $\tau$=0.2 the cavitation region is clearly visible only at $U$=40 m/s and 45 m/s. At $\tau$=0.6 the cavitation region is visible at all speeds, however at $U$=40 m/s and 45 m/s it has already reached the trailing edge of the specimen, whereas at $U$=30 m/s and at $U$=35 m/s it has expanded less significantly.
(a) $U$=30 m/s $\tau$ =1
(b) $U$=35 m/s $\tau$ =1
(c) $U$=40 m/s $\tau$ =1
(d) $U$=45 m/s $\tau$ =1
@@
By looking at the signals recorded by the probes located at the rear, all of them exhibit a descent to the vapour pressure value. The start time of the descent is consistent across all speeds, while the descent itself becomes sharper as the horizontal speed increases. At $U$=40 m/s and 45 m/s, all probes, except for P15 and P17 (located on $r4$), display a rise towards the atmospheric value due, to the occurrence of ventilation, which is consistent with what shown by the underwater images. The rise starts earlier at higher speed. The probes P15 and P17, being located in the region where the longitudinal curvature is rather high, remains in the cavitation region and are not reached by the ventilation front.

### v2

@@ 2.5 Test Conditions
The test conditions that are discussed in the present paper are listed in Table 2. The guide inclination angle is $\beta=\arctan(V/U)$. For all tests examined $V/U=1.5/40=0.0375$, so $\beta=2.1476^{\circ}$ (see again Figure 4). It is worth noticing that, due to the use of bungee ropes to accelerate the trolley, a very precise control of the speed cannot be achieved, which is the reason why the actual values of the horizontal speed are not exactly the nominal expected values, also shown in the table. It should be noted that in the discussion from now on, however, reference will be made to the nominal velocity values and not to the actual ones reported in the table.
Water temperature was measured on a near-daily basis using a thermometer. Recorded values ranged from 15 to 18 C. The air temperature was approximately 20 C.
@@ 3 Results and discussion
3.1 Definition of the impact phase
For the purpose of the present study, the data analysis is limited to the impact phase, which lasts from the time of first contact of the specimen with the water surface, $t_{0}$, up to the time $t_{E}$ at which the root of the forward propagating spray reaches the leading edge of the specimen. As done in Iafrati and Grizzi (2019), $t_{0}$ can be approximated as the time at which the first pressure probe exhibits a pressure growth, whereas $t_{E}$ is estimated by linear extrapolation starting from the times at which the pressure peak passes over the location of the probes P28 and P30 (Iafrati, 2016). As shown in Figure 4, for the shapes S2 and S3 at a pitch angle of 6 the estimation of $t_{0}$ based on the time history of the pressure measured by the probe P17 is expected to be quite accurate, being the probe P17 located very close to the point of first contact H. Instead, for the shape S1B the first contact point is located between P21 and P24, therefore the estimate of $t_{0}$ is more uncertain. The estimate based on pressure measurements is corrected by computing the time duration of the impact phase from the underwater videos and subtracting it from $t_{E}$.
It is important to note that tests conducted on specimens can effectively capture the hydrodynamic behaviour of the entire fuselage ditching (with the imposed trajectory, horizontal speed, and constant pitch angle), as long as the wetted area during water entry remains within the specimen’s boundaries. This definition of the impact phase ensures that the forward part of the wetted area stays within the specimen’s area until it reaches the leading edge. Since the front and rear walls are located above the specimen’s profile, they do not affect the pressures underneath, at least until the impact phase ends. Minor discrepancies may occur, after the cavitation region or the wetted area reaches the trailing edge or the lateral edges, but these differences are believed to be small.
@@ 3.2 Hydrodynamics phenomena during the water entry
Qualitative considerations on the hydrodynamics of the water entry of the double curvature specimens can be made through a comparative analysis of the pressure time histories and the underwater high-speed videos. Such an analysis is conducted on the test of the shape S2, advancing with a horizontal velocity UU=40 m/s, velocity ratio V/UV/U=0.0375 and pitch angle $\alpha$= 6. Video frames taken at different $\tau$, displaying the flow evolution on the bottom of the specimen, are shown in Figure 7.
@@
Figure 7: Frames of the underwater video at different $\tau$ during the water entry of the shape S2 at a nominal horizontal speed of UU=40 m/s and at a pitch angle of $\alpha$=6. The specimen moves from left to right.
@@
(a) Probe positions: C – separation between zero-curvature part on the right and finite-curvature part on the left; H – first contact point at $\alpha$ = 6.
@@
Figure 8: Pressure probe time histories row by row for the shape S2 moving at UU=40 m/s, at a pitch angle of $\alpha$=6.
From the underwater frames shown in Figure 7 it is possible to recognize the wetted area as the darker region. During the early stage of impact, i.e. in Figure 7(b,c,d)(b,c,d), the wetted area grows in all directions. The brighter zone around the wetted area represents the spray region.
@@
By focusing on the longitudinal plane (midplane) in the trolley-fixed reference frame, the results of the two-dimensional case of an asymmetrical wedge entering water with a horizontal velocity (Judge et al., 2004; Xu et al., 2008; Krastev et al., 2018) can provide some elements for the interpretation of the flow phenomena in the present case. In fact, as a consequence of the kinematic asymmetry of the flow due to the horizontal speed, the pressure on the front part of the specimen is higher than that at the rear. From a theoretical standpoint, it is expected that a velocity singularity will occur at the apex (Judge et al., 2004). In real flows, a strong local velocity acceleration is observed there, accompanied by a pressure drop, which is similar to what occurs on the suction side of hydrofoils (Faltinsen, 2005; Newman, 1971). The flow field presents also similarities with that of the vertical water entry of inclined asymmetric wedges (Semenov and Iafrati, 2006; Riccardi and Iafrati, 2004). In Xu et al. (2008), in which a wedge entering water with both a vertical and a horizontal velocity is studied using a self-similar approach, even if the velocity ratio $U/V$ is much lower than the present case, a local pressure drop to negative values is observed at the wedge apex.
An alternative interpretation of the pressure behaviour can be obtained by focusing on the flow in the transverse plane, in an earth-fixed reference frame, by exploiting the so called 2D+$t$ approach (Fontaine and Tulin, 1998; Iafrati and Broglia, 2010). In the front part the plate always penetrates into the water, whereas in the rear part, depending on the local inclination of the tangent to the body surface in the longitudinal plane and on the $U/V$ velocity ratio, the fluid flow problem changes from a water entry to a water exit one (Tassin et al., 2013) and when this happens, negative pressures develop. The development of low and even negative pressures occurs very shortly after the first contact with water and, based on the above considerations, the pressure drop is expected to be higher if the gradient in the slope of the body surface is more significant, which means in presence of larger longitudinal curvature.
@@
As observed in Iafrati and Grizzi (2019), the pressure drop can be so high that the vapour pressure value may be reached. With a water temperature between 15 and 18  C, the vapour pressure relative to the atmospheric value, denoted as $p_{\textrm{vap}}$, is in the range -99.6 and-99.2 kPa. This explain the formation of a cavitation region (Acosta, 1973; Wang et al., 2001; Brennen, 2014), as shown in the underwater images of Figure 7 at the non-dimensional time $\tau\geq 0.2$. In the trolley-fixed reference frame, the cavitation region has a sort of reverse ”D” shape, which follows the pressure distribution on the bottom of the specimen. As the specimen enters in water, the forward front of the cavitation region maintains rather straight and located about the point of curvature change C, whereas the rear edge of the cavitation region spreads backwards towards the trailing edge of the specimen. Once the cavitation region reaches the trailing edge of the specimen, the water vapour inside it, at vapour pressure, comes into contact with air at atmospheric pressure. As a consequence, air is suddenly entrained into the water vapour cavity, forming a ventilation zone that expands in the opposite direction, i.e towards the leading edge. This behaviour is typical of the so called cavitation-ventilation regime, described in more detail in Section 3.3 based on the analysis of the water entry of the specimen S3 (Iafrati and Grizzi, 2019; Iafrati et al., 2020) and also occurs for S2 at the same horizontal speed.
@@
The expansion of the cavitation region can be also recognized by observing the time histories of the pressure probes, see Figure 8. The passage of the rear edge of the cavitation region corresponds to the abrupt drops to the vapour pressure value. Moving backwards, i.e. from row $r4$ to row $r1$, the instant at which the drops occur for the midline probes (P17, P13, P9 and P4) increases. This time delay is related to the propagation velocity of the rear edge of the cavitation region. Comparing the pressure time histories across the same rows, it can be noted that pressure drops of the midline probes are delayed with respect to those of the probes at the sides, due to the combined effect of of the body shape and of the curvature of the rear edge of the cavitation region. When ventilation starts and air is entrained within the cavity, pressures start increasing progressively, reaching the zero-pressure value. It is worth noting that, in contrast to the pressure drop associated with the propagation of the cavitation region, the rise in pressure values during this phase is gradual and occurs at extremely close time instants for all pressure probes positioned on the same row. The increase in pressure values occurs later in time, when progressing from row $r1$ to $r4$. This clearly indicates that the ventilation front is advancing at an extremely high speed, as previously observed for the shape S3 in Iafrati and Grizzi (2019).
@@
A comprehensive global overview of the shape and evolution of the cavitation region can be derived by constructing pressure iso-lines using the available pressure data from row $r1$ to $r5$ at different time instants. In Figure 9, iso-lines are superimposed on the corresponding video frames at $\tau=0.3$, $\tau=0.6$, and from $\tau=0.7$ to $\tau=0.8$ with a finer non-dimensional time resolution of $\Delta\tau=0.02$.
(a) $\tau$=0.30
(b) $\tau$=0.60
(c) $\tau$=0.70
(d) $\tau$=0.72
(e) $\tau$=0.74
(f) $\tau$=0.76
(g) $\tau$=0.78
(h) $\tau$=0.8
@@
Figure 9: Pressure iso-lines in the rear part of the specimen overlapped to the underwater video frames for the test of shape S2 at U=40 m/s, at $\alpha=6^{\circ}$ at the specfied $\tau$ values. The symmetry about the mid-line is exploited.
@@
Observing the pressure time histories, it can be observed that ventilation occurs within the time interval between $\tau\approx 0.7$ and 0.8. By focusing the pressure iso-lines over this time interval, as shown in Figure 9, it is possible to gain more insight on the ventilation dynamics. In fact, it is observed that the pressure begins to rise back to the atmospheric value first at the centre, since the cavitation region reaches the trailing edge first there. However, the ventilation front develops rapidly and at about $\tau$=0.78. becomes roughly parallel to the transverse direction, as also observed for the shape S3 in Iafrati and Grizzi (2019).
@@
Figure 10: Time histories of the forces recorded by the front (red) and rear (blue) load cells and of the total normal load (black), for shape S2 at 6 pitch angle and UU=40 m/s. The vertical red and blue lines indicate the start of cavitation and of ventilation respectively. The solid and dashed lines refer to the first and second test repeat respectively (denoted as R1 and R2).
@@
3.3 Effect of the ditching speed
@@
As previously noted in Iafrati and Grizzi (2019); Iafrati et al. (2020), significant alterations in the hydrodynamics of the water entry are observed when the horizontal speed is varied for the shape S3. As shown in Table2, when the horizontal speed is varied while keeping the ratio $V/U$ constant, the vertical speed also varies. In particular, at least four regimes were identified: a no-cavitation regime (for $U$ lower than about 27 m/s), in which no cavitation is observed; an incipient cavitation regime (between $U\approx$ 27 m/s and 35 m/s) in which a cavitation region expanding in time only up to a certain $\tau$ can be seen; an intermediate regime, (between $U\approx$ 35 m/s and 37 m/s) (denoted in Iafrati and Grizzi (2019) as Full Cavitation Bubble and Cavitation Alternate with Ventilation), in which the cavitation region reaches the trailing edge, followed by the formation of a local ventilation region, which does not expand in relevant way; finally a cavitation-ventilation regime ($U$ higher than 37 m/s), in which the ventilation region exhibits a successive fast expansion towards the leading edge of the specimen.
@@
In the present work, the effect of the horizontal speed is analysed for the shape S2, entering water with a pitch angle of 6 and at velocity ratio of $V/U=0.0375$. The analysis is performed in terms of pressures and loads and through the observation of the underwater images taken at two non-dimensional times, i.e. $\tau=0.2$ and $\tau=0.6$. A comparison is presented between the time histories of pressures recorded by the probes located on the midline and on line $l3$. (Figure 6). In order to make the comparison easier, the probes lying along rows r5r8r5-r8 and r1r4r1-r4 are plotted in separate graphs. The results for UU=30 m/s and 35 m/s are shown in Figure 11, whereas those for UU=40 m/s and 45 m/s in Figure 12.
(a) UU=30 m/s - Frame $\tau$ =0.2
(b) UU=35 m/s - Frame $\tau$ =0.2
(c) UU=30 m/s - Frame $\tau$ =0.6
(d) UU=35 m/s - Frame $\tau$ =0.6
(e) $p$ front probes - midline - UU= 30 m/s
(f) $p$ front - probes midline - UU= 35 m/s
(g) $p$ front probes - line l3 UU= 30 m/s
(h) $p$ front probes - line l3 UU= 35 m/s
(i) $p$ midline - rear probes - UU= 30 m/s
(j) $p$ midline - rear probes - UU= 35 m/s
(k) $p$ line l3 UU= 30 m/s
(l) $p$ line l3 UU= 35 m/s
@@
Figure 11: Front and rear pressure probe time histories and under-water video frames comparison for the tests of shape S2, at a pitch angle 6 and UU=30 m/s and 35 m/s.
(a) UU=40 m/s - Frame $\tau$ =0.2
(b) UU=45 m/s - Frame $\tau$ =0.2
(c) UU=40 m/s - Frame $\tau$ =0.6
(d) UU=45 m/s - Frame $\tau$ =0.6
(e) $p$ front probes - midline - UU= 40 m/s
(f) $p$ front - probes midline - UU= 45 m/s
(g) $p$ front probes - line l3 UU=40 m/s
(h) $p$ front probes - line l3 UU=45 m/s
(i) $p$ midline - rear probes - UU=40 m/s
(j) $p$ midline - rear probes - UU=45 m/s
(k) $p$ line l3 UU=40 m/s
(l) $p$ line l3 UU= 45 m/s
@@
Figure 12: Front and rear pressure probe time histories and under-water video frames comparison for the tests of shape S2, at a pitch angle 6 and UU=40 m/s and 45 m/s.
@@
As shown in Figures 11 and 12 the probes located in the front, i.e. ahead of the point H, are characterized by the typical sharp rise, followed by a gentle decay. It is worth noting that, for all probes, the non-dimensional time $\tau$ at which the pressure peak is attained is almost independent of the horizontal speed, whereas the peak value increases with the speed. As expected, the peaks recorded by the probes along the line $l3$ occur at a later time with respect to the peaks recorded by the probes at the midline in the same row. This is a consequence of the combined effect of the transverse curvature of the specimen and of the curvature of the spray root, which is associated with the possibility for the water to escape from the sides (Iafrati, 2016).
@@
The analysis of the pressure measurements of the probes located behind the point H is less straightforward and the support of the underwater images is needed for a correct interpretation. In fact, looking at the underwater frames in Figures 11 and 12 it can be seen that at $\tau$=0.2 the cavitation region is clearly visible only at UU=40 m/s and 45 m/s. At $\tau$=0.6 the cavitation region is visible at all speeds, however at UU=40 m/s and 45 m/s it has already reached the trailing edge of the specimen, whereas at UU=30 m/s and at UU=35 m/s it has expanded less significantly.
(a) UU=30 m/s $\tau$ =1
(b) UU=35 m/s $\tau$ =1
(c) UU=40 m/s $\tau$ =1
(d) UU=45 m/s $\tau$ =1
@@
By looking at the signals recorded by the probes located at the rear, all of them exhibit a descent to the vapour pressure value. It should be noted that at U=30 m/s, possibly due to the incipient cavitation regime, signals from rear probes P2, P4, P7, P9, and P11 exhibit oscillatory behaviour during this descent. The start time of the descent is consistent across all speeds, while the descent itself becomes sharper as the horizontal speed increases. At UU=40 m/s and 45 m/s, all probes, except for P15 and P17 (located on $r4$), display a rise towards the atmospheric value due, to the occurrence of ventilation, which is consistent with what shown by the underwater images. The rise starts earlier at higher speed. The probes P15 and P17, being located in the region where the longitudinal curvature is rather high, remains in the cavitation region and are not reached by the ventilation front.